What is the diameter of a 1.00-m length of tungsten wire whose resistance is 0.10 Ω ?
Answer #1

Length of the wire, L= 1.0m

Resistance of wire, R= 0.10 \Omega

Resistivity of tungsten, ρ= 5.6 × 10-812m

Let "d" is the diameter of the wire and "r" is the radius of the wire, then we can write Area of wire as

4

Now using the formula of resistance,

R= \frac{\rho L}{A}

using all given values in above,

  0.1= \frac{5.6 \times 10^{-8} \times 1.0 }{\pi \times \frac{d^2}{4}}

  d^2= \frac{5.6 \times 10^{-8} \times 1.0 }{3.14 \times 0.1 \times 0.25}

d= \sqrt{0.7133 \times 10^{-6}}=0.844 \times 10^{-3}m= 0.844mm(ANS)